OpenStudy (anonymous):

Let f(x) = 10x^2. Find f(3 + h).

OpenStudy (nenadmatematika):

=10(3+h)^2=10(9+6h+h^2)=10h^2+60h+90

OpenStudy (anonymous):

oh dang, I got that, I just put it in my homework the wrong way, Thanks for the help though

OpenStudy (nenadmatematika):

you're welcome :D

OpenStudy (anonymous):

what if you are looking for f ' (3) you have f(3+h) = 10h^2 +60h +90 and 10h +60 and you have f(3+h) - f(3)/h